# How Post-Training Reshapes LLMs: A Mechanistic View on Knowledge, Truthfulness, Refusal, and Confidence

### v2

@@ Abstract
Post-training is essential for the success of large language models (LLMs), transforming pre-trained base models into more useful and aligned post-trained models. While plenty of works have studied post-training algorithms and evaluated post-training models by their outputs, it remains understudied how post-training reshapes LLMs internally. In this paper, we compare base and post-trained LLMs mechanistically from four perspectives to better understand post-training effects. Our findings across model families and datasets reveal that: (1) Post-training does not change the factual knowledge storage locations, and it adapts knowledge representations from the base model while developing new knowledge representations; (2) Both truthfulness and refusal can be represented by vectors in the hidden representation space. The truthfulness direction is highly similar between the base and post-trained model, and it is effectively transferable for interventions; (3) The refusal direction is different between the base and post-trained models, and it shows limited forward transferability; (4) Differences in confidence between the base and post-trained models cannot be attributed to entropy neurons. Our study provides insights into the fundamental mechanisms preserved and altered during post-training, facilitates downstream tasks like model steering, and could potentially benefit future research in interpretability and LLM post-training. Our code is publicly available at HZD01/post-training-mechanistic-analysis.
@@ 1 Introduction
The success of large language models (LLMs) has standardized a training paradigm consisting of pre-training and post-training. Post-training transforms a pre-trained base model into more useful and aligned post-trained models (Grattafiori et al., 2024; OpenAI, 2023; Jiang et al., 2023; Lambert et al., 2024, inter alia). Initially introduced to improve instruction-following capabilities (Ouyang et al., 2022; Wei et al., 2022), post-training has evolved to serve versatile purposes, including but not limited to making models more truthful (Lin et al., 2022; OpenAI, 2023; Lambert et al., 2024), safety alignment by enabling models to refuse harmful instructions (Bai et al., 2022; Grattafiori et al., 2024), and calibrating the model’s output confidence (OpenAI, 2023).
@@
Research on post-training has predominantly focused on algorithms such as Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) and improving LLMs’ ability in downstream tasks such as reasoning (Kumar et al., 2025) and math (Liu et al., 2024b). These studies mainly treat the LLM as a black box, and only evaluate its outputs externally (Zhou et al., 2023; Wen et al., 2024). However, it remains unclear how post-training affects the mechanisms of LLMs and whether the model is fundamentally altered internally. Such a mechanistic understanding can help us better use post-trained LLMs and potentially design better post-training methods.
@@
Recent research studies have started to examine the mechanistic effect of post-training and reveal interesting findings. However, this direction is still underexplored, given these efforts are still algorithm-centric (Lee et al., 2024), model-specific (Panickssery et al., 2024), task-format-specific (Panickssery et al., 2024), or rely on learning an extra model like Sparse Autoencoders (SAEs) on top of the LLM instead of direct analysis (Kissane et al., 2024b).
@@
In this work, we systematically and mechanistically study the post-trained (post) model, on top of the pre-trained (base) model. We compare the base and post models internally from four perspectives: knowledge storage and representation, internal belief of truthfulness, refusal behavior, and confidence. These perspectives represent fundamental capabilities that determine an LLM’s real-world utility and safety. post models are expected to preserve knowledge learned during pre-training, improve truthfulness, enhance refusal of harmful inputs, and show a different level of confidence from the base model. We specifically focus on two post model types: a final model that went through all post-training stages, commonly called the instruct model, and a model with only supervised fine-tuning on top of base, commonly called the sft model. While some other perspectives, such as reasoning and instruction-following, are also important, they involve complex, multi-step processes that are not well-captured by current interpretability tools. Therefore, our work focuses on four perspectives that can be rigorously measured and mechanistically interpreted, providing a solid foundation for understanding internal mechanisms during post-training.
@@
Figure 1: Summary of our analysis. (a) Knowledge: knowledge-storage locations are overlapping between base and post models as the difference is small; (b) Truthfulness: the truthfulness direction is similar between base and post models; (c) Refusal: the refusal direction is different between base and post models; (d) Confidence: the difference in confidence between base and post models cannot be attributed to entropy neurons as they are largely overlapping (numbers are entropy neurons’ IDs).
@@
For each perspective, we choose the most suitable tool from the LLM interpretability toolbox. For the first perspective, we utilize the widely used knowledge locating technique, causal tracing (Meng et al., 2022), to investigate the storage and representation of knowledge. We discover that locations for storing the same knowledge in base and post models are similar, and post model adapts the original knowledge representations while developing new ones. For the second perspective of truthfulness, based on previous papers’ discovery that the concept of truthfulness can be represented as a direction in the hidden representation space (Marks & Tegmark, 2024; Li et al., 2024; Panickssery et al., 2024; Bürger et al., 2024), we learn a vector linearly representing truthfulness in the model’s hidden space, referred to as the “truthfulness direction”. For the two directions learned for base and post models, we find that they have high cosine similarity and can be effectively transferred for truthfulness intervention. For the third perspective, we learn a “refusal direction” similar to the truthfulness direction in the hidden representation space (Arditi et al., 2024). We find that the transferability of such refusal direction is only effective backward (from post to base) but not forward (from base to post). Last, we analyze the confidence of base and post models through the lens of entropy neurons, which contributes to forming the LLM’s output confidence (Stolfo et al., 2024; Gurnee et al., 2024). Our analysis reveals that entropy neurons of base and post models have similar distributions, leading us to the conclusion that these neurons do not contribute significantly to the observed confidence differences between the base and post models. We illustrate our main conclusions in Figure 1.
@@
Our analysis from the four perspectives reveals both the kept and the altered internal mechanisms by post-training, which could benefit future research and applications in interpretability and LLM post-training. Given some internal mechanisms are mostly developed during pre-training and not significantly altered by post-training, such as factual knowledge and the truthfulness direction. We can leverage the transferability to develop for example truthfulness-oriented procedures on the base model and apply it to the post model conveniently. For the mechanisms that are altered or developed during post-training, such as refusing harmful instructions, there are also possibilities to efficiently improve base’s ability by applying the backward transfer from post.
@@ 2 Related Work
Mechanistic Interpretability (MI) of Post-training MI aims to understand internal mechanisms of models (Elhage et al., 2021; Wang et al., 2022; Templeton et al., 2023; Nanda et al., 2023, inter alia). Recently, a growing body of research starts to analyze LLM post-training through the MI lens. Lee et al. (2024) studied how DPO unlearns toxicity in LLM, finding that rather than removing toxic-promoting vectors, the model learns distributed offsets to bypass them. Panickssery et al. (2024) discovered that Llama-2 base and instruct models have similar steering vectors for answering multiple choice questions. Kissane et al. (2024a) showed that refusal directions can be transferred from instruct models to base models. Kissane et al. (2024b) revealed that the SAEs trained on the base model can reconstruct the activations of the instruct model. However, these investigations do not directly and generally reveal the post-training effect, whereas we do a comprehensive study of different models and datasets and investigate post-training’s effect from four critical perspectives.
@@
Knowledge Storage and Representation Geva et al. (2021) showed that transformer MLP layers function as key-value memories, with keys corresponding to input representations and values inducing output distributions. Dai et al. (2022) identified specific “knowledge neurons” in MLPs that encode facts. To detect knowledge-storage locations and edit them, Meng et al. (2022) introduced causal tracing (activation patching) and edited knowledge through targeted weight changes. These studies show that knowledge in LLMs can be localized and modified through causal intervention techniques. In this work, we use a variant of causal tracing to study the effect of post-training on knowledge storage.
@@
Internal Belief of Truthfulness Recent research demonstrates that LLMs encode the belief of truthfulness linearly in their representation space as a “truthfulness direction”. Azaria & Mitchell (2023) identified truthfulness signals in model activations, while Burns et al. (2024) developed unsupervised methods to extract these signals using logical consistency. Li et al. (2024) leveraged truthfulness directions to improve truthfulness through activation steering. Later, Marks & Tegmark (2024) introduced the mass-mean (MM) probe. Similarly, Panickssery et al. (2024) uses difference-in-means to identify the direction by computing the difference between mean activation vectors of true and false statements. Additionally, Bürger et al. (2024) discovered a universal two-dimensional truthfulness subspace across various LLMs, and Liu et al. (2024a) showed that training the direction on more datasets makes it more robust, suggesting that a universal truthfulness hyperplane may exist. We employ MM probe (Marks & Tegmark, 2024) and show that the truthfulness direction persists after post-training.
@@
Refusal Behavior Refusing to answer harmful instructions is a key objective of post-training. Recent research has revealed that this behavior is linearly mediated by a vector as a “refusal direction” (Arditi et al., 2024). This direction can be used to undermine the model’s ability to refuse harmful requests. Similarly, research on prompt-driven safeguarding has shown that safety prompts typically move input queries in the refusal direction in the representation space  (Zheng et al., 2024). Further research has shown this direction can also be learned on base models, or transferred from an instruct model to a base model (Kissane et al., 2024a). Our work extends the study to a more systematic comparison of the refusal direction learned on base and different post models across model families.
@@
Confidence and Entropy Neurons Confidence calibration is another key objective of post-training. Studies have shown that post-trained models tend to be less calibrated, with instruct models being overconfident compared to base models (Tian et al., 2023). One line of research is to understand LLM’s confidence with verbalized output (Tian et al., 2023; Xiong et al., 2024), using prompting and sampling strategies to generate multiple responses and compute consistency. Another line of work analyzes confidence to show that specialized neurons within LLMs regulate uncertainty (Katz & Belinkov, 2023; Gurnee et al., 2024; Stolfo et al., 2024). Among them, Gurnee et al. (2024) discovered “entropy neurons” that have high weight norms but minimal direct logit effects. They modulate uncertainty by influencing layer normalization to scale down logits. Our work examines the changes in entropy neurons after post-training to understand its effect on confidence.
3 Notations and Experiments Settings
@@
Notations Throughout the paper, we denote layers as $l\in[L]$ and token positions as $i\in[I]$, where $L$ is the layer number and $I$ is the input length. We use notations like $\mathcal{D}_{\text{harmless}}^{\text{train}}$ for datasets, with superscript for train/test, and subscript for the dataset’s type. The representation at layer $l$ and position $i$ of an input statement $s$ is denoted as $h_{i}^{l}(s)$. We use $\mathbf{W_{U}}\in\mathbb{R}^{|\mathcal{V}|\times d_{model}}$ for the unembedding matrix, with vocabulary $\mathcal{V}$, and $\mathbf{W}_{\text{out}}$ for the output weights vector of a given neuron in the last-layer MLP.
@@
Models We mainly conduct experiments on two representative LLM model families: Llama-3.1-8B/Instruct (Grattafiori et al., 2024) and Mistral-7B-v0.3/Instruct (Jiang et al., 2023). The original model release does not include sft models of this size, so we use widely recognized external sft models: Llama-3.1-Tulu-3-8B-SFT, which finetunes Llama-3.1-8B on the tulu-3-sft-mixture dataset (Lambert et al., 2024), and Mistral-7B-Base-SFT-Tulu2 (Feuer et al., 2025), which finetunes Mistral-7B-v0.3 on the tulu-v2-sft-mixture dataset (Ivison et al., 2023). For refusal experiments, we additionally include Qwen-1.5-0.5B/Instruct (Bai et al., 2023) and Gemma-2-9B/Instruct (Team et al., 2024) following experiment settings in Arditi et al. (2024). For confidence experiments, we additionally include Llama-2-7B/Instruct models (Touvron et al., 2023) following Stolfo et al. (2024). To further demonstrate that our findings could generalize to different models sizes, especially larger models, we include experiments on Llama-2-13B/Instruct (Touvron et al., 2023) models for all perspectives in Appendix F.
@@
Datasets For the knowledge and truthfulness perspectives, we use datasets from (Marks & Tegmark, 2024; Bürger et al., 2024; Azaria & Mitchell, 2023), where each sub-dataset contains simple and unambiguous statements that are either true or false from diverse topics. For example, cities contains statements about cities and their countries, following the format “The city of [city] is in [country]”. The unambiguity and clear dataset make it easy to analyze LLMs. To eliminate the concern that the datasets might be out-of-distribution for post-training, we curate a dataset that is in-distribution for sft models. We curate the tulu_extracted dataset from the tulu-3-sft-mixture dataset (Lambert et al., 2024), which was used to finetune the Llama-3.1-8B-SFT model. We ensure every statement from tulu_extracted also appears in the tulu-v2-sft-mixture dataset (Ivison et al., 2023), so it is also in-distribution for the Mistral-SFT model. For experiments on the refusal perspective, we follow Arditi et al. (2024) to use advbench (Zou et al., 2023) for harmful inputs and alpaca (Taori et al., 2023) for harmless inputs. Dataset details are explained in Appendix A.

### v3

@@ Abstract
Post-training is essential for the success of large language models (LLMs), transforming pre-trained base models into more useful and aligned post-trained models. While plenty of works have studied post-training algorithms and evaluated post-training models by their outputs, it remains understudied how post-training reshapes LLMs internally. In this paper, we compare base and post-trained LLMs mechanistically from four perspectives to better understand post-training effects. Our findings across model families and datasets reveal that: (1) Post-training does not change the factual knowledge storage locations, and it adapts knowledge representations from the base model while developing new knowledge representations; (2) Both truthfulness and refusal can be represented by vectors in the hidden representation space. The truthfulness direction is highly similar between the base and post-trained model, and it is effectively transferable for interventions; (3) The refusal direction is different between the base and post-trained models, and it shows limited forward transferability; (4) Differences in confidence between the base and post-trained models cannot be attributed to entropy neurons. Our study provides insights into the fundamental mechanisms preserved and altered during post-training, facilitates downstream tasks like model steering and benefits future research in interpretability and LLM post-training. Our code is publicly available at HZD01/post-training-mechanistic-analysis.
@@ 1 Introduction
The success of large language models (LLMs) has standardized a training paradigm consisting of pre-training and post-training. Post-training transforms a pre-trained base model into more useful and aligned post-trained models (grattafiori2024llama; openai2024gpt4technicalreport; jiang2023mistral7b; Tulu, inter alia). Initially introduced to improve instruction-following capabilities (training_LM_to_follow_instructions; wei2022finetunedlanguagemodelszeroshot), post-training has evolved to serve versatile purposes, including but not limited to making models more truthful (lin2022truthfulqameasuringmodelsmimic; openai2024gpt4technicalreport; Tulu), safety alignment by enabling models to refuse harmful instructions (bai2022constitutionalaiharmlessnessai; grattafiori2024llama), and calibrating the model’s output confidence (openai2024gpt4technicalreport).
@@
Research on post-training has predominantly focused on algorithms such as Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage) and Reinforcement Learning from Human Feedback (RLHF) (RLHF) and improving LLMs’ ability in downstream tasks such as reasoning (LLM_Post_Training_Reasoning; acemath). These studies mainly treat the LLM as a black box, and only evaluate its outputs externally (zhou2023instructionfollowingevaluationlargelanguage; wen2024benchmarkingcomplexinstructionfollowingmultiple). However, it remains unclear how post-training affects the mechanisms of LLMs and whether the model is fundamentally altered internally. Such a mechanistic understanding can help us better use post-trained LLMs and potentially design better post-training methods.
@@
Recent research studies have started to examine the mechanistic effect of post-training and reveal interesting findings. However, this direction is still underexplored, given these efforts are still algorithm-centric (lee2024mechanisticunderstandingalignmentalgorithms), model-specific (panickssery2024steeringllama2contrastive), task-format-specific (panickssery2024steeringllama2contrastive), or rely on learning an extra model like Sparse Autoencoders (SAEs) on top of the LLM instead of direct analysis (sae_finetuning).
@@
In this work, we systematically and mechanistically study the post-trained (post) model, on top of the pre-trained (base) model. We specifically focus on two post model types: a model that went through all post-training stages, commonly called the instruct model, and a model with only supervised fine-tuning on top of base, commonly called the sft model. We compare the base and post models internally from four perspectives: knowledge storage and representation, internal belief of truthfulness, refusal behavior, and confidence. These perspectives represent fundamental capabilities that determine an LLM’s real-world utility and safety. post models are expected to preserve knowledge learned during pre-training, improve truthfulness, enhance refusal of harmful inputs, and show a different level of confidence from the base model. While some other perspectives, such as reasoning and instruction-following, are also important, they involve complex, multi-step processes that are not well-captured by current mechanistic interpretability tools. Therefore, our work focuses on those four perspectives above that can be rigorously measured and mechanistically interpreted, providing a solid foundation for understanding the internal mechanism updates during post-training.
@@
Figure 1: Summary of our analysis and findings. (a) Knowledge: A difference heatmap showing base and post models have negligible location differences for storing the same knowledge; (b) Truthfulness: A PCA plot showing the truthfulness directions are similar in base and post models; (c) Refusal: A PCA plot showing the refusal directions of base and post models are quite different; (d) Confidence: A Venn diagram of entropy neuron IDs showing the difference in confidence between base and post models cannot be fully attributed to entropy neurons as they largely overlap.
@@
For each perspective, we choose the most suitable tool from the LLM interpretability toolbox for analysis, and we illustrate our main findings in Figure 1. For the first perspective, we adopt the widely used knowledge locating technique, causal tracing (meng2022locating), to investigate the storage and representation of knowledge. We discover that locations for storing the same knowledge in base and post models are similar, and post model adapts the base knowledge representations while developing new ones. For the second perspective of truthfulness, based on the previous discovery that truthfulness can be represented as a direction in the model’s hidden representation space (marks2024geometrytruthemergentlinear; li2024inferencetimeinterventionelicitingtruthful; panickssery2024steeringllama2contrastive; bürger2024truthuniversalrobustdetection), we learn a linear vector representing truthfulness, referred to as the “truthfulness direction.” For the two directions learned for base and post models, we find that they have high cosine similarity and can be effectively transferred for truthfulness intervention. For the third perspective, we learn a “refusal direction” similar to the truthfulness direction in the hidden representation space (arditi2024refusallanguagemodelsmediated). We find that the transferability of such refusal direction is only effective backward (from post to base) but not forward (from base to post). Lastly, we analyze the confidence of base and post models through the lens of entropy neurons, which contribute to the confidence of the LLM’s output (stolfo2024confidenceregulationneuronslanguage; gurnee2024universalneuronsgpt2language). Our analysis reveals that entropy neurons of base and post models have similar distributions, leading us to the conclusion that these neurons are not the determining factor of the observed confidence differences between the base and post models.
@@
Our analysis from the four perspectives reveals both the kept and the altered internal mechanisms by post-training, which could benefit future research and applications in interpretability and LLM post-training. As our results suggested, some internal mechanisms are mostly developed during pre-training and not significantly altered by post-training, such as factual knowledge storage and the truthfulness direction. We can thus leverage their transferability to develop procedures on the base model and apply them to the post model conveniently, for example, for truthfulness steering. For the mechanisms that are altered or developed during post-training, such as refusing harmful instructions, it is also possible to efficiently improve base’s ability by applying the backward transfer from post.
@@ 2 Related Work
Mechanistic interpretability of post-training Mechanistic interpretability aims to understand internal mechanisms of models (elhage2021mathematical; wang2022interpretabilitywildcircuitindirect; templeton2023monosemantic; nanda2023progressmeasuresgrokkingmechanistic, inter alia). Recently, a growing body of research starts to analyze LLM post-training through the Mechanistic interpretability lens. lee2024mechanisticunderstandingalignmentalgorithms studied how DPO unlearns toxicity in LLM, finding that rather than removing toxic-promoting vectors, the model learns distributed offsets to bypass them. panickssery2024steeringllama2contrastive discovered that Llama-2 base and instruct models have similar steering vectors for answering multiple choice questions. BaseLLMsRefuseToo showed that refusal directions can be transferred from instruct models to base models. sae_finetuning revealed that the SAEs trained on the base model can reconstruct the activations of the instruct model. However, these investigations do not directly and generally reveal the post-training effect, whereas we do a comprehensive study of different models and datasets and investigate post-training’s effect from four critical perspectives.
@@
Knowledge storage and representation geva2021transformerfeedforwardlayerskeyvalue showed that transformer MLP layers function as key-value memories, with keys corresponding to input representations and values inducing output distributions. dai2022knowledgeneuronspretrainedtransformers identified specific “knowledge neurons” in MLPs that encode facts. To detect knowledge-storage locations and edit them, meng2022locating introduced causal tracing (activation patching) and edited knowledge through targeted weight changes. These studies show that knowledge in LLMs can be localized and modified through causal intervention techniques. In this work, we use a variant of causal tracing to study the effect of post-training on knowledge storage.
@@
Internal belief of truthfulness Recent research demonstrates that LLMs encode the belief of truthfulness linearly in their representation space as a “truthfulness direction”. azaria2023internalstatellmknows identified truthfulness signals in model activations, while burns2024discoveringlatentknowledgelanguage developed unsupervised methods to extract these signals using logical consistency. li2024inferencetimeinterventionelicitingtruthful leveraged truthfulness directions to improve truthfulness through activation steering. Later, marks2024geometrytruthemergentlinear introduced the mass-mean probe. Similarly, panickssery2024steeringllama2contrastive uses difference-in-means to identify the direction by computing the difference between mean activation vectors of true and false statements. Additionally, bürger2024truthuniversalrobustdetection discovered a universal two-dimensional truthfulness subspace across various LLMs, and liu2024universaltruthfulnesshyperplaneinside showed that training the direction on more datasets makes it more robust, suggesting that a universal truthfulness hyperplane may exist. We employ mass-mean probe (marks2024geometrytruthemergentlinear) and show that the truthfulness direction persists after post-training.
@@
Refusal behavior Refusing to answer harmful instructions is a key objective of post-training. Recent research has revealed that this behavior is linearly mediated by a vector as a “refusal direction” (arditi2024refusallanguagemodelsmediated). This direction can be used to undermine the model’s ability to refuse harmful requests. Similarly, research on prompt-driven safeguarding has shown that safety prompts typically move input queries in the refusal direction in the representation space  (zheng2024promptdrivensafeguardinglargelanguage). Further research has shown this direction can also be learned on base models, or transferred from an instruct model to a base model (BaseLLMsRefuseToo). Our work extends the study to a more systematic comparison of the refusal direction learned on base and different post models across model families.
@@
Confidence and entropy neurons Confidence calibration is another key objective of post-training. Studies have shown that post-trained models tend to be less calibrated, with instruct models being overconfident compared to base models (tian2023justaskcalibrationstrategies). One line of research is to understand LLM’s confidence with verbalized output (tian2023justaskcalibrationstrategies; xiong2024llmsexpressuncertaintyempirical), using prompting and sampling strategies to generate multiple responses and compute consistency. Another line of work analyzes confidence to show that specialized neurons within LLMs regulate uncertainty (katz2023visitvisualizinginterpretingsemantic; gurnee2024universalneuronsgpt2language; stolfo2024confidenceregulationneuronslanguage). Among them, gurnee2024universalneuronsgpt2language discovered “entropy neurons” that have high weight norms but minimal direct logit effects. They modulate uncertainty by influencing layer normalization to scale down logits. Our work examines the changes in entropy neurons after post-training to understand its effect on confidence.
3 Notations and Experimental Settings
@@
Notations Throughout the paper, we denote layers as $l\in[L]$ and token positions as $i\in[I]$, where $L$ is the number of model layers and $I$ is the input length. We use notations like $\mathcal{D}_{\text{harmless}}^{\text{train}}$ for datasets, with superscript for train/test subset, and subscript for the dataset’s type, which might be omitted if the context is clear. The representation at layer $l$ and position $i$ of an input statement $s$ is denoted as $h_{i}^{l}(s)$. We use $\mathbf{W_{U}}\in\mathbb{R}^{|\mathcal{V}|\times d_{model}}$ for the unembedding matrix, and $\mathbf{w}_{\text{out}}\in\mathbb{R}^{d_{model}}$ for the output weights of a given neuron in the last-layer MLP, with $\mathcal{V}$ stands for the vocabulary, and $d_{model}$ for the model’s hidden space dimension.
@@
Models We mainly conduct experiments on two representative families of LLMs: Llama-3.1-8B/Instruct (grattafiori2024llama) and Mistral-7B-v0.3/Instruct (jiang2023mistral7b). The original model releases do not include corresponding sft models, so we use widely recognized external sft models: Llama-3.1-Tulu-3-8B-SFT, which finetunes Llama-3.1-8B on the tulu-3-sft-mixture dataset (Tulu), and Mistral-7B-Base-SFT-Tulu2 (feuer2025styleoutweighssubstancefailure), which finetunes Mistral-7B-v0.3 on the tulu-v2-sft-mixture dataset (ivison2023camelschangingclimateenhancing). For refusal experiments, we additionally include Qwen-1.5-0.5B/Instruct (bai2023qwen) and Gemma-2-9B/Instruct (team2024gemma) following experiment settings in arditi2024refusallanguagemodelsmediated. For confidence experiments, we additionally include Llama-2-7B/Instruct models (touvron2023llama) following stolfo2024confidenceregulationneuronslanguage. To further demonstrate that our findings could generalize to different model sizes, especially larger models, we perform experiments on Llama-2-13B/Instruct (touvron2023llama) models for all perspectives.
@@
Datasets For the knowledge and truthfulness perspectives, we start with a group of datasets from marks2024geometrytruthemergentlinear and curate them to fit specific experiments we run. Each of these datasets contains simple and unambiguous statements from diverse topics that are either true or false. For example, the dataset cities contains statements about cities and their countries, following the format “The city of [city] is in [country]”. While these datasets are independent of post-training, we also perform our analysis on datasets that are actually used for post-training to reveal the post-training effect on in-distribution data. Specifically, we construct the tulu_extracted dataset by sampling factual statements from the tulu-3-sft-mixture dataset (Tulu), which was used to finetune Llama-3.1-8B to Llama-3.1-8B-SFT. For each extracted statement, we generate a counterfactual counterpart to form true–false pairs. We ensure that all sampled statements also appear in the tulu-v2-sft-mixture dataset (ivison2023camelschangingclimateenhancing), making tulu_extracted in-distribution for both Llama-3.1-8B-SFT and Mistral-7B-SFT. For experiments on the refusal perspective, we follow arditi2024refusallanguagemodelsmediated to use advbench (zou2023universal) for harmful inputs and alpaca (alpaca) for harmless inputs. Dataset details are explained in Appendix A.